OpenStudy (anonymous):

MEDAL GIVEN if you can solve problem: if log(base 3)x^2 - log(base 3) x=2 what is X?

OpenStudy (anonymous):

Apply power rule of logarithms

OpenStudy (anonymous):

i applied the rule. still can't figure it out.

OpenStudy (anonymous):

Actually combine the two logs using the quotient rule.

OpenStudy (anonymous):

What does that give you?

OpenStudy (anonymous):

i did the power rule and then combined (divided them) so i have 2 log base 3 x over log base 3 x =2

OpenStudy (anonymous):

log (base 3) of x^2 - log (base 3) of x =2 2 log (base 3) of x - log (base 3) of x =2 log (base 3) of x = 2 After this just rearrange from log form to exponential form.

OpenStudy (anonymous):

no you can't do that...if you bring the 2 out front then you can't combine the two logs...leave the power on there and just combine the logs ... what do you get?

OpenStudy (anonymous):

My bad what cloudy did is good too

OpenStudy (anonymous):

thank youuuuuuuuu

OpenStudy (anonymous):

With the other guy's method: log (base 3) x^2 - log (base 3) x = 2 log (base 3) ((x^2)/x) =2